### v1

# Investigating Twin Star Equation of States in Light of Recent Astrophysical Observations

Abstract
Twin stars are predicted to exist in nature if the hadron-to-quark phase transition is strong enough to form a new branch of hybrid stars, separated from the branch of neutron stars. It is essential to study the hybrid equations of state that can form twin stars, with a primary focus on estimating the onset of phase transition and its strength. We adopt an agnostic approach, using transition energy density, transition pressure, the discontinuity strength, and a constant speed of sound for quark matter as our parameter space to construct a large possibility of hybrid equations of state, and thereby encapsulating a comprehensive picture of the twin star scenario. First we report the complete conditions on our parameter space imposed by the general relativistic hydrostatic equilibrium solutions. For a fixed transition energy density and speed of sound for quark matter, we define distinct ranges of transition pressures based on the allowed strengths of discontinuity. Below a maximum transition pressure, a range of discontinuity exists that increases as the transition pressure decreases. Thereby, we identify the loci of the limits on discontinuities as the ‘witch-hat’ curves. Based on the causality limit, the witch-hat curves can be punctured or incomplete. Strong constraints on this picture are drawn from the inferences from GW170817 and the NICER measurements, especially from the recent measurements of PSR J0614–3329 and PSR J0437–4715, which point towards a softer equation of state around the canonical mass of neutron stars. We computed the maximum mass for twin stars to be $2.05~M_{\odot}$, the allowed strongest discontinuity in rest-mass density to be $7.76\rho_{\mathrm{sat}}$, and the upper bound on transition rest-mass density to be $4.03\rho_{\mathrm{sat}}$. Subsequently, we compute the implications of the stiffness of the quark matter equation of state on this picture. Different confidence levels for observational inferences are considered to assess the extent of inclusion (and rejection) of hybrid equations of state and, consequently, their effects on the limits of the maximum mass of twin stars, the maximum transition density and the allowed strongest discontinuity.
I Introduction
The pursuit of understanding the equation of state (EoS) of matter inside a Neutron Star (NS) is a traditional one [1, 2, 3]. It remains an important revisit as we continue to observe more of the universe with increasingly advanced technologies, particularly by measuring these compact objects and witnessing the violent events involving them. NSs possess strong gravity such that the matter at the core is compressed beyond nuclear saturation density $\rho_{\mathrm{sat}}\simeq 2.7\times 10^{14}~\mathrm{g/cm}^{3}$ [4]. The unavailability of such circumstances on earth-based environments makes NSs the natural laboratories in the universe to study the behaviour of matter at low temperatures and high densities [5, 6, 7, 8]. At such extreme conditions, a phase transition (PT) from hadronic matter (HM) to quark matter (QM) is expected from Quantum Chromodynamics (QCD) [9].
The objective to achieve an integrated picture of QCD has been approached from various directions. In the principles of Chiral Effective Field Theory [10, 11, 12, 13, 14], the EoS of the cold matter is well-anticipated to be in hadronic state upto $\sim 1.1\rho_{\mathrm{sat}}$ [15, 16, 17, 18, 19, 20, 21, 22]. From the earth-based experiments involving the relativistic heavy-ion collision has mainly probed the EoS of matter at high temperatures and low densities [23, 24, 25, 26, 27]. Theoretical framework of lattice QCD [28, 29, 30] efficiently probes the matter at low densities, however, interfered by sign problem [31, 32] when extended to densities at the regimes of NSs. Perturbative QCD (pQCD) calculations [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44] have computed on behaviour of matter asymptotically high densities ($\sim 40\rho_{\mathrm{sat}}$).
To probe the density regime that resides inside the core of an NS ($1\sim 10~\rho_{\mathrm{sat}}$), there are a couple of directions to move forward. Strong constraints have been drawn from the precise mass-radius measurements made by the NICER X-ray observations [45, 46, 47, 48, 49, 50, 51, 52, 53]. Importantly, the recent mass-radius measurement of PSR J0614–3329 [52] and PSR J0437–4715 [53] point towards a softer EoS, in contrast to the earlier measurements of J0030+0451 [45, 46, 47] and J0740+6620 [48, 49, 50, 51], which suggested otherwise. Additional assistance came from the multi-messenger astronomy, mainly the observation of gravitational waves from binary merger of NSs [54], which has inferred constraints on the tidal deformability of NSs [55, 56, 57, 58, 59, 60], which also steers the investigation to a softer composition of matter around the canonical mass of NS ($1.4~M_{\odot}$). On the other side, the mass measurement of massive pulsars ($\gtrsim 2.0~M_{\odot}$) [61, 62, 63, 64, 65] advocate for a relatively stiffer EoS.
Despite the progress made, there can still exist a plethora of thermodynamically consistent EoSs that comply with the observational status discussed above [66, 67, 68, 69, 70]. The attempts to probe the properties of PT remain important yet inadequate. Primarily, the onset of PT is not known a priori, and the nature of first-order PT is proposed by models that include—($i$) Maxwell construction–the PT is modelled as a transition using discontinuity jump in the density at a constant pressure, ($ii$) Gibbs construction–the HM and QM coexist in a mixed phase inside a region of pressure-density [71, 72]. With such prospects and uncertainties in hand, model-agnostic approaches are useful for encapsulating all possibilities in an EoS describing the behaviour of matter inside NSs [73, 70, 74, 75, 76].
A typical hybrid EoS with Maxwell-type PT constructed out of an agnostic approach consists of a pure hadronic part at lower densities and a pure quark part at higher densities separated by a PT region described by a discontinuous jump in the density at constant transition pressure. Under the general relativistic hydrostatic equilibrium conditions governed by the Tolman Oppenheimer Volkoff (TOV) equations [77, 78], such EoSs indicate the existence of hybrid stars (HSs), which contain a quark seed within a hadronic shell. Interestingly, within the constraints imposed by astrophysical observations, the discontinuity in the density jump that describes the strength of the PT can be large enough to give rise to a third branch of HSs, separated from the second branch of NSs [79, 80]. In this context, the first branch of compact objects refers to the family of white dwarfs. Such scenarios form the twin stars (TSs) that reside on the third branch (smaller radius, more compact) but have equal mass with respect to NSs on the second branch (larger radius, less compact).
TSs are the direct manifestation of strong hadron-to-quark PT. Observation of such candidates of compact objects can confirm the occurrence of PT in the astrophysical scenarios and severely constrain the EoS [81, 66, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94], especially if detections are made in the high-mass regime [95, 96, 97, 98]. On the other hand, the formation of TSs in realistic scenarios and their dynamical evolution remain a serious pursuit [99, 100]. Several studies have presented the possibility of realistic EoSs that can construct such species [101, 102, 103, 81, 104, 105, 106, 107, 108, 109, 110, 111, 112]. A few works have studied the oscillation modes [113, 114, 115, 116, 117], QCD trace anomaly [118], rotation [119, 120, 121, 122], magnetic fields [123, 124] and temperature effects [125, 126, 127] on these compact objects.
En route to uncovering the TS scenario, it is essential to sketch a comprehensive (if not the complete) picture of TSs, considering the large possibility of hybrid EoSs that can coexist with recent observational constraints. In this study, we adopt an agnostic approach to build an ensemble of EoSs that can capture TS solutions. In particular, we allow transition energy density and transition pressure to vary freely to allow a larger set of hadronic EoSs within the causality limit to form hybrid EoSs. Systematic analysis of such an inclusive set of EoSs reveals more of the TS scenario presented in the previous literature. We revisit the criteria for TS solutions as given by the TOV equations to gain a comprehensive picture beyond the existing literature. In particular, we define the ‘witch-hat’ curves for a choice of transition energy density and stiffness of quark matter EoS. The effects of causality and the properties of PT on these curves are exhaustively addressed, revealing specific conditions that can give rise to punctured or incomplete witch-hat curves. Previous studies have constructed witch-hat like curves only for a limited or fixed hadronic EoS [83, 80, 79]. After establishing the framework under hydrostatic equilibrium within the limits of causality, we constrain the picture using recent observational inferences and set bounds on the properties of PT and the maximum mass of TSs. In particular, the recent inferences from PSR J0614–3329 and PSR J0437–4715 are rather strong and constitute an important constraint to be included in new studies.
The article is organized in the following way. In Sec. II, we discuss the agnostic approach adopted to construct the ensemble of hybrid EoS, define our parameter space, and outline our strategy for locating the third branch. In Sec. III, we discuss our results in three main parts—approach from TOV, constraints from observation and possible bounds on the TS scenario. Finally, we conclude our work with remarks in Sec. IV.
Figure 1: [Left] Diagram indicating the different parts of EoS, that are controlled by different parameters—transition energy density ($e_{\mathrm{tr}}$), transition pressure ($P_{\mathrm{tr}}$), the strength of discontinuity in energy density ($\Delta e$) and constant speed of sound for quark matter EoS ($c^{2}_{\mathrm{s,QM}}$). [Right] Schematic diagram of MMRR sequence of an EoS that constructs TSs. The second (purple) and third (orange) stable branches are separated by a sequence of unstable branch (black dashed curve). The maximum mass on the second and third branches ($M_{\mathrm{TOV,2}}$ and $M_{\mathrm{TOV,3}}$) is indicated by purple and red stars, respectively. The sequence of unstable stars begins beyond these stellar configurations. The minimum mass on the third branch is indicated by a red circle. The twin region is defined as the mass range in which equal-mass stellar models are present on both stable branches.
II Formalism
II.1 Equation of State
We employ an agnostic approach to systematically construct EoSs that provide a macroscopic description of matter properties, particularly the connection between the thermodynamic quantities—rest-mass density ($\rho$), energy density ($e$), and pressure ($P$). The EoSs are computed using the following parameters:
  • transition energy density ($e_{\mathrm{tr}}$)
  • transition pressure ($P_{\mathrm{tr}}$)
  • discontinuity in energy density ($\Delta e$)
  • constant speed of sound for quark EoS ($c^{2}_{\mathrm{s,QM}}$)
The independent choices of $e_{\mathrm{tr}}$ and $P_{\mathrm{tr}}$ decide the stiffness of hadronic EoS. These parameters are sufficient to describe hybrid EoSs with first-order PT, which mainly involves large discontinuities. We recall that TSs have been categorised based on the maximum mass of the second branch ($M_{\mathrm{TOV,2}}$) and the third branch ($M_{\mathrm{TOV,3}}$) in Ref. [80], which are as follows:
  • Category I: $M_{\mathrm{TOV,2}}\geq 2.0~M_{\odot}$ and $M_{\mathrm{TOV,3}}\geq 2.0~M_{\odot}$, particularly associated with higher $e_{\mathrm{tr}}$.
  • Category II: $M_{\mathrm{TOV,2}}\geq 2.0~M_{\odot}$ and $M_{\mathrm{TOV,3}}\leq 2.0~M_{\odot}$, particularly associated with larger $\Delta e$.
  • Category III: $M_{\mathrm{TOV,2}}\leq 2.0~M_{\odot}$ and $M_{\mathrm{TOV,3}}\geq 2.0~M_{\odot}$.
  • Category IV: $M_{\mathrm{TOV,2}}\leq 1.0~M_{\odot}$ and $M_{\mathrm{TOV,3}}\leq 2.0~M_{\odot}$, particularly associated with low $P_{\mathrm{tr}}$.
See Fig. 3 in Ref. [83] for schematic reference. Each EoS is set up using the four sections explained below:
  • Crust Region: At smaller densities ($\lesssim 0.4\rho_{\mathrm{sat}}$), a fixed crust EoS is built using four piecewise polytropes, which are taken from Table II in Ref. [74]. The addition of this part remains common throughout all the EoSs constructed for this study.
  • Hadronic Region ($e<e_{\mathrm{tr}}$): An additional polytrope is added above the crust EoS to mimic the hadronic matter, with adiabatic index $\gamma$ as a free parameter. The thermodynamic quantities—$P$ and $e$ are given in terms of $\rho$ as,
    $P(\rho)=k\rho^{\gamma},\quad e(\rho)=\rho(a+1)+\dfrac{P}{\gamma-1}$ (1)
    The polytropic constant ($k$) and $a$ are computed using the boundary condition to ensure continuity of $P(\rho)$ and $e(\rho)$ between the crust and the hadronic EoS. The independent choices for $e_{\mathrm{tr}}$ and $P_{\mathrm{tr}}$ allow for finding an appropriate value of $\gamma$ and transition rest-mass density ($\rho_{\mathrm{tr}}$). These quantities are numerically computed using the equations,
    $\rho_{\mathrm{tr}}=\left(\dfrac{P_{\mathrm{tr}}}{k}\right)^{{1}/{\gamma}},\quad e_{\mathrm{tr}}=\rho_{\mathrm{tr}}(1+a)+\dfrac{P_{\mathrm{tr}}}{\gamma-1}$ (2)
    For a fixed $e_{\mathrm{tr}}$, a higher (lower) value of $P_{\mathrm{tr}}$ will set a higher (lower) value of $\gamma$, resulting in a stiffer (softer) hadronic EoS.
  • Phase Transition Region ($e_{\mathrm{tr}}<e<e_{\mathrm{tr}}+\Delta e$): The PT region is computed after fixing $e_{\mathrm{tr}}$ and $P_{\mathrm{tr}}$, allowing a free choice to be made for $\Delta e$. This value decides the strength of the discontinuity of PT. The pressure will remain constant in this region as given by $P_{\mathrm{tr}}$. This step leads to the construction of Maxwell-type PT [72]. Since the chemical potential remains constant in the region of PT, the $\Delta\rho$ can be computed as,
    $\Delta\rho=\dfrac{\rho_{\mathrm{tr}}\Delta e}{P_{\mathrm{tr}}+e_{\mathrm{tr}}}$ (3)
  • Quark Region ($e>e_{\mathrm{tr}}+\Delta e$): The quark part is modelled using the constant speed of sound (CSS) parametrisation with $c_{\mathrm{s,QM}}^{2}$ [79, 80]. The pressure in the quark region is given as,
    $P(e)=P_{\mathrm{tr}}+c_{\mathrm{s,QM}}^{2}e$ (4)
Fig 1 (left) shows the schematic diagram of the EoS construction method described above. First, we fix a value of $e_{\mathrm{tr}}$, after which the $P_{\mathrm{tr}}$ can be varied from an upper bound set by the limit of causality, that is, the maximum value of $P_{\mathrm{tr}}$ is set at a value where $c^{2}_{\mathrm{s}}(P_{\mathrm{tr}},e_{\mathrm{tr}})=1$. However, if the central pressure of $M_{\mathrm{TOV}}$ of hadronic EoS is smaller than this upper limit of pressure, then the maximum value of $P_{\mathrm{tr}}$ is set to the central pressure of the stellar configuration at $M_{\mathrm{TOV}}$. This rule assures that we induce PT in the relevant positions of an MMRR sequence and not in the cores of unstable stellar models beyond $M_{\mathrm{TOV}}$.
On the other hand, the lower limit of $P_{\mathrm{tr}}$ is set to a value where such a central pressure would construct an NS of $1~M_{\odot}$, which means $M_{\mathrm{TOV,2}}>1~M_{\odot}$. However, if an unstable branch is formed by an EoS that constructs $M_{\mathrm{TOV,2}}=1~M_{\odot}$, we only increase $\Delta e$ up to a limit where the third branch forms a minimum mass up to $0.7~M_{\odot}$. We do not construct EoSs which form a minimum mass below $0.7~M_{\odot}$ in the third branch. These choices are made to restrict our search to reasonable PT densities. In theory, such EoS (with large $\Delta e$ and very small $P_{\mathrm{tr}}$) can be constructed where TSs can form at very low masses ($\lesssim 0.5~M_{\odot}$), which are far from the NS range (minimum mass of NSs is known to be $\sim 1.1~M_{\odot}$ [128, 129]). Searches for such TS solutions are avoided in this study. Whereby, we do not survey the TSs of category IV, which are not responsible for constructing massive TSs or assisting in setting upper limits on the $e_{\mathrm{tr}}$ and $\Delta e$. Moreover, small $P_{\mathrm{tr}}$ at large $e_{\mathrm{tr}}$ will make extremely soft hadronic EoSs, resulting in hybrid EoSs that will be unable to support massive NSs ($\gtrsim 2.0~M_{\odot}$). Hence, we find TSs of category IV will be strictly ruled out at large $e_{\mathrm{tr}}$. At lower $e_{\mathrm{tr}}$, the acceptance of such TSs can be anticipated from our analysis and does not require specific exploration. See Refs. [80, 102, 103, 107, 111], which have especially examined TSs of category IV.
For the PT region, the minimum $e_{\mathrm{tr}}$ ($\rho_{\mathrm{tr}}$) used in this study is $3\times 10^{14}~\mathrm{g/cm}^{3}$ ($\approx 1\rho_{\mathrm{sat}}$). The CET sufficiently predicts that matter resides as hadronic matter up to $\sim 1.1\rho_{\mathrm{sat}}$ [130, 131]. We construct a bunch of hadronic EoSs by uniformly varying $P_{\mathrm{tr}}$ in the described range and set the quark matter EoS on the other side by fixing a value for $c^{2}_{\mathrm{s,QM}}$. The $\Delta e$ (and hence $\Delta\rho$) is adaptively varied to find the limits on its allowed range, such that the given pair of hadronic and quark EoS can construct TS solutions. The accuracy of determining the limits on $\Delta e$ ($\Delta\rho$) is considered up to $10^{12}~\mathrm{g/cm}^{3}$ ($\approx 10^{-2}\rho_{\mathrm{sat}}$).
Figure 2: Allowed region of EoS space with fixed $e_{\mathrm{tr}}$ that allows TS solutions. The grey region indicates the range of hadronic EoS considered in constructing the hybrid EoSs. The solution-less region (red) indicates the range of $P_{\mathrm{tr}}$ inside which TSs do not form. The left panel shows the MMRR sequences of two EoSs in this region with different $\Delta e$. The bound region (blue) is separated from the red region using $P^{\mathrm{max}}_{\mathrm{tr}}$ at which a $\Delta e$ exists that allows TS solutions. The middle panels show that the EoS with $\Delta e$ within the allowed values of $\Delta e$ constructs TSs. The sequence of the TSs is indicated by cyan in the inset. The unbound region (green) region only contains the minimum value of $\Delta e$ above which TS solutions are possible. The right panel shows that the EoS with large value of $\Delta e$ form TSs. The sequence of TSs is indicated by turquoise in the inset. The loci of minimum and maximum $\Delta e$ are indicated in black, which we identify as the ‘witch-hat’ curve. The black dashed line is an extrapolation of the maximum $\Delta e$.
II.2 Locating third branch
Fig. 1 (right) shows the schematic diagram of the MMRR sequence that has separated second and third branches containing TSs. To detect the TS solutions in the third branch, we first locate the central density of the maximum mass configuration on the second branch (purple curve), which is constructed from the hadronic EoS. Based on the Seidov criteria [132], upon increasing the central density to build further stellar models, the EoS may allow for two scenarios. The immediate sequence of HSs may continue to be in the second branch, or they may fall into the unstable branch (discussed as scenarios (b) and (d) in Fig. 2 of reference [79]). In either of the situations, the $M_{\mathrm{TOV,2}}$ is computed as indicated with a purple star in Fig. 1 (right). Consequently, an unstable branch (black dashed curve) is recognised only if the mass of the subsequent stellar models decreases up to $\Delta M=10^{-3}~M_{\odot}$ with respect to the purple star. The quark EoS can be stiff enough that a third branch (orange curve) may appear after a sequence of unstable HSs. The third branch is detected only if the mass of the subsequent stellar models increases up to $\Delta M=10^{-3}~M_{\odot}$ from the minimum mass in the unstable branch (indicated with a red circle). Finally, the third branch reaches its $M_{\mathrm{TOV,3}}$ (indicated with a red star), and the subsequent stellar models form another unstable branch (black dashed curve). Once we numerically obtain this sequence of branches within the tolerance $\Delta M$, the EoS is recognised to have produced TS solutions.
The detection margin of $\Delta M=10^{-3}~M_{\odot}$ is suitable for our study, as this accuracy is several orders of magnitude above the observational accuracy [133, 134, 135, 52, 53, 45, 49, 51, 63, 61, 62, 64, 65, 136, 137]. Besides, detection margins $\Delta M<10^{-3}~M_{\odot}$ lead to ambiguity between the physical stable/unstable branches versus a local fluctuation due to numerical calculation, which are highly probable when TOV equations have to handle EoSs with strong and discontinuous PT.
III Results
We discuss the results from two main perspectives. Within the described parameter space, we analyse the conditions on allowed EoSs that give rise to the TS scenario, purely governed by TOV solutions. Based on the established theoretical framework, we apply recent observational constraints to further develop inferences, including upper bounds on transition densities, the strength of the PT discontinuity, and the maximum mass of TSs.
III.1 Inferences from TOV solutions
Fig. 2 shows a schematic picture of the region in EoS space with a fixed $e_{\mathrm{tr}}$, indicating the allowed region of $\Delta e$ that can form TSs. The grey region indicates the range of hadronic EoSs constructed according to the fixed $e_{\mathrm{tr}}$ and the allowed range of $P_{\mathrm{tr}}$ set by our formalism. By varying the $P_{\mathrm{tr}}$ in this range and fixing a value for $c^{2}_{\mathrm{s,QM}}$, we identify three distinct regions based on the allowed conditions on $\Delta e$, which are explained below.
Solution-less Region (Red)—In this region, TSs do not form for any value of $\Delta e$, as shown in Fig. 2 (left panel). The inset shows the MMRR sequences of selected EoSs that have $P_{\mathrm{tr}}$ in this region, emphasising that a third branch does not appear in these sequences after entering the unstable branch beyond $M_{\mathrm{TOV,2}}$. For such combinations of $e_{\mathrm{tr}}$ and $P_{\mathrm{tr}}$, the quark EoS is not stiff enough to form another stable branch beyond the unstable branch.
Bound Region (Blue)—The maximum value of $P_{\mathrm{tr}}$ in this region (indicated as $P^{\mathrm{max}}_{\mathrm{tr}}$) is extracted numerically, at which TSs can form for a unique value of $\Delta e$. For $P_{\mathrm{tr}}<P^{\mathrm{max}}_{\mathrm{tr}}$, there exists a range of $\Delta e$, in which TSs can form, as shown in Fig. 2 (middle panel). We selected three EoSs in this region, two of which lie outside the allowed region of $\Delta e$. The inset showing the respective MMRR sequences indicates that only the EoS having $\Delta e$ within the allowed region forms a third branch beyond the second branch, which is separated by an unstable branch. The sequence of TSs is indicated by the cyan curve on the third branch. Decreasing $P_{\mathrm{tr}}$ results in an increase in the maximum limit of $\Delta e$ and thereby increasing the range of $\Delta e$ allowed for TS solutions. At the lowest value of $P_{\mathrm{tr}}$ in this region, the maximum limit of $\Delta e$ is no more found in our numerical setup. However, an extrapolation using the locus of maximum $\Delta e$ in this region indicates the growing (rather blowing up) nature of maximum $\Delta e$ for the smaller values of $P_{\mathrm{tr}}$.
Unbound Region (Green)—This region contains only a minimum value of $\Delta e$, as there is no maximum limit on $\Delta e$. As shown in Fig. 2 (right), we selected two EoSs from either side of the minimum $\Delta e$. The respective MMRR sequences show that TSs (indicated by the turquoise curve on the third branch) can form for much larger values of $\Delta e$. The plot shows the end of this region at a minimum value of $P_{\mathrm{tr}}$, where $M_{\mathrm{TOV,2}}=1~M_{\odot}$. However, that is only a restriction applied to the EoS survey in this study. The formation of TSs is still possible below the minimum value of $P_{\mathrm{tr}}$ accounted in this study. That is, the locus of minimum $\Delta e$ will continue to exist below the limit of this region.
Figure 3: The effects of change in $e_{\mathrm{tr}}$ and $c^{2}_{\mathrm{s,QM}}$ on witch-hat curve. [Left] The vertical fading lines indicate the $e_{\mathrm{tr}}$. The respective Seidov lines of each $e_{\mathrm{tr}}$ are indicated in grey solid lines. [Right] The Seidov lines align in the modified axes.
In this article, we identify the loci of minimum and maximum $\Delta e$ with varying $P_{\mathrm{tr}}$ and fixed $e_{\mathrm{tr}}$ as the ‘witch-hat’ curve (because the set of loci resembles a witch’s hat) to simplify the discussion around this picture. We note that a witch-hat curve can be computed for a given combination of the parameters—($e_{\mathrm{tr}}$, $c^{2}_{\mathrm{s,QM}}$). A point under the witch-hat curve represents a unique EoS (informing the values of $P_{\mathrm{tr}}$ and $\Delta e$ in addition to $e_{\mathrm{tr}}$ and $c^{2}_{\mathrm{s,QM}}$) that constructs TS solutions. Similar curves have been discussed before in the literature, but with a fixed hadronic EoS, which constrains $P_{\mathrm{tr}}$ given the choice of $e_{\mathrm{tr}}$ [79, 80]. While treating $P_{\mathrm{tr}}$ and $e_{\mathrm{tr}}$ as free variables, we are able to explore a larger possibility of EoS, resulting in new features about the witch-hat curve that have not been discussed previously.
In Fig. 3 (left), we show that the witch-hat curves heavily depend on the variation in ($e_{\mathrm{tr}}$, $c^{2}_{\mathrm{s,QM}}$). We find that an increase in $e_{\mathrm{tr}}$ leads to an increment in the size of the witch-hat curve, indicating a larger region is allowed for $P_{\mathrm{tr}}$ and $\Delta e$. For a fixed $e_{\mathrm{tr}}$, the lower values of $c^{2}_{\mathrm{s,QM}}$ form a smaller witch-hat with a lower value of $P^{\mathrm{max}}_{\mathrm{tr}}$. Interestingly, we observe a puncture on the tip of the witch-hat curve formed by ($e_{\mathrm{tr,b}}$, $c^{2}_{\mathrm{s,QM1}}$). Here, the loci of the minimum and maximum of $\Delta e$ do not meet, which is different from the expected picture discussed in Fig. 2. We find that for such a case, the $P^{\mathrm{max}}_{\mathrm{tr}}$ would be beyond the causality limit, which is also the upper limit for the survey range of $P_{\mathrm{tr}}$. If we were to allow a survey of hadronic EoSs beyond the causality limit, the witch-hat curve would have been completed. As we lower the value of $e_{\mathrm{tr}}$, for instance ($e_{\mathrm{tr,a}}$, $c^{2}_{\mathrm{s,QM1}}$), we find that the witch-hat curve is even more incomplete, where we only find the locus of minimum $\Delta e$. Such scenarios result in $P_{\mathrm{tr}}$ at the causality limit to be present in the unbound region (green) of Fig. 2, and hence the incomplete witch-hat curve would not have a well-defined bound region (blue). Hence, for the incomplete witch-hat curves, we define $P^{\mathrm{max}}_{\mathrm{tr}}$ at the causality limit. However, for smaller values of $c^{2}_{\mathrm{s,QM}}$, the witch-hat curve can be complete. For example, the witch-hat curve for ($e_{\mathrm{tr,b}}$, $c^{2}_{\mathrm{s,QM2}}$) is complete, whereas the witch-hat curve for ($e_{\mathrm{tr,b}}$, $c^{2}_{\mathrm{s,QM1}}$) is punctured. This effect is expected from the fact that smaller $c^{2}_{\mathrm{s,QM}}$ forms the witch-hat curve in the lower range of $P_{\mathrm{tr}}$.
In Fig. 3 (right), we compute the witch-hat curves in $\Delta e/e_{\mathrm{tr}}-P_{\mathrm{tr}}/e_{\mathrm{tr}}$ axis plot, which has been majorly used to study this subject in the past literature [79, 80]. Since the Seidov lines align in this plot, unlike in Fig. 3 (left), the effect of $e_{\mathrm{tr}}$ and $c^{2}_{\mathrm{s,QM}}$ on witch-hat curves can be compared directly. We demonstrate that the picture surrounding the witch-hat curves discussed in this study agrees with the special cases presented in earlier studies. Therefore, the bifurcation of the region under the witch-hat curve based on categories of TSs, as defined in Ref. [80] (see Fig. 6 therein), remains consistent and applicable in our work.
Figure 4: Allowed regions under the witch-hat curve for three different values of $e_{\mathrm{tr}}$, constrained from observational measurements ($2\sigma$ confidence). The vertical fading lines indicate the value of $e_{\mathrm{tr}}$. For all the EoSs, the quark part is described with $c^{2}_{\mathrm{s,QM}}=1.0$. The inset shows the MMRR sequences of the EoSs with ($P_{\mathrm{tr}}$, $\Delta e$) marked by circles inside the allowed region. Only the stable branches of MMRR sequences are shown, where the second (third) branch is indicated using light (dark) colour in the respective colour theme used for the $e_{\mathrm{tr}}$. Black dashed lines are marked at $1.0~M_{\odot}$ and $2.0~M_{\odot}$ for the identification of TS categories.
Figure 5: Emphasis on particular observational measurements ($2\sigma$ confidence) that are responsible for constraining various regions under the witch-hat curve for three different values of $e_{\mathrm{tr}}$. The vertical fading lines indicate the value of $e_{\mathrm{tr}}$. The quark EoS description is set with $c^{2}_{\mathrm{s,QM}}=1.0$. The inset shows the MMRR sequences, along with the particular observational constraint that they violate. These MMRR sequences are constructed with the EoSs with ($P_{\mathrm{tr}}$, $\Delta e$) marked by circles outside the allowed region. Only the stable branches of MMRR sequences are shown, where the second (third) branch is indicated using light (dark) colour in the respective colour theme used for the $e_{\mathrm{tr}}$.
Figure 6: Change in the picture of allowed regions of $\Delta e$ with varying $c^{2}_{\mathrm{s,QM}}$ (0.33, 0.55, 0.77 and 1.0) along with imposition of the constraints from observational measurements at different confidences ($1\sigma$, $2\sigma$ and $3\sigma$). The choice of $c^{2}_{\mathrm{s,QM}}$ remains the same over a column, and the choice of confidence remains constant over a row. The background contour is taken from [70], indicating that the range of allowed EoSs in this study lies over (however, not completely within) the constrained region provided in the literature. The vertical fading lines indicate the value of $e_{\mathrm{tr}}$.
We observed that the TOV solutions can put only an upper limit on $P_{\mathrm{tr}}$ (if not by the causality limit). No upper limit is provided for the $e_{\mathrm{tr}}$ or $\Delta e$ (at small values of $P_{\mathrm{tr}}$). Interestingly, we find that the current conglomerate of the observational measurements can provide strong constraints on this picture, including an upper bound for $e_{\mathrm{tr}}$ and appropriate limits on $\Delta e$.
III.2 Constraints from recent observational measurements
For developing further constraints from astrophysical point of view, we incorporated observational inference from NICER mass-radius measurements (PSR J0614–3329 [52], PSR J0437–4715 [53], PSR J0030+0451 [45, 46, 47], and PSR J0740+6620 [48, 49, 50, 51]), the mass measurement of massive pulsar PSR J0348+0432 [63], and the tidal deformability from GW170817 [60]. Fig. 4 shows that the allowed region under a witch-hat curve can be heavily constrained using these observational constraints (using $2\sigma$ confidence), and only selective regions of ($P_{\mathrm{tr}}$, $\Delta e$) are allowed for each value of $e_{\mathrm{tr}}$. We present this picture using three specific values of $e_{\mathrm{tr}}$ (with $c^{2}_{\mathrm{s,QM}}=1.0$), which sufficiently explain the constraints received from observational measurements at different regimes of transition densities:
  • small: $e_{\mathrm{tr}}=3\times 10^{14}~\mathrm{g/cm}^{3}$, $\rho_{\mathrm{tr}}\approx 1\rho_{\mathrm{sat}}$
  • intermediate: $e_{\mathrm{tr}}=6\times 10^{14}~\mathrm{g/cm}^{3}$, $\rho_{\mathrm{tr}}\approx 2\rho_{\mathrm{sat}}$
  • large: $e_{\mathrm{tr}}=9\times 10^{14}~\mathrm{g/cm}^{3}$, $\rho_{\mathrm{tr}}\approx 3\rho_{\mathrm{sat}}$
In Fig. 5, we emphasise that different observational measurements become relevant to constrain different regions under the witch-hat curves, and that different combinations of observational measurements come into play at different regimes of $e_{\mathrm{tr}}$. Firstly, we show that, for all three values of $e_{\mathrm{tr}}$, the maximum values of $\Delta e$ are constrained by the current observational inferences (see more discussion in Sec. III.3). In particular, GW170817, PSR J0437–4715 and PSR J0614–3329 assist in constraining the respective maximum $\Delta e$ at small, intermediate and large values of $e_{\mathrm{tr}}$. Since larger values of $\Delta e$ result in more compact third branches (smaller radii), the maximum $\Delta e$ is bound to be constrained by the observational measurements at some point, as all of them have a minimum value in the uncertainty of the radii. However, it is subject to the fact that we did not find an EoS using our formalism that forms a second branch satisfying all the constraints by itself and then forms a third branch in the regime of the unbound region (green) discussed in Fig. 2. In such a case, finding a maximum limit on $\Delta e$ using observational constraints would not be possible. The maximum values of $\Delta e$ ($\Delta\rho$) were noted to be $1.59\times 10^{15}~\mathrm{g/cm}^{3}$ ($4.71\rho_{\mathrm{sat}}$), $1.77\times 10^{15}~\mathrm{g/cm}^{3}$ ($4.98\rho_{\mathrm{sat}}$) and $2.94\times 10^{15}~\mathrm{g/cm}^{3}$ ($7.76\rho_{\mathrm{sat}}$) at small, intermediate and large $e_{\mathrm{tr}}$, respectively.
Secondly, in Fig. 4, we observe that, even though the larger $e_{\mathrm{tr}}$ results in larger witch-hat curves, however, observational measurements tend to constrain these regions more tightly. We find that the collection of allowed (satisfying causality) hadronic EoSs at larger $e_{\mathrm{tr}}$ is softer, resulting in lower $M_{\mathrm{TOV}}$. As a result, at extremely large values of $e_{\mathrm{tr}}$, the hybrid EoSs consisting of such hadronic parts won’t be stiff enough to satisfy the mass constraint from PSR J0348+0432. Hence, there exists a maximum value of $e_{\mathrm{tr}}$ for the TS solution in the light of current observational constraints (see more discussion in Sec. III.3).
For small value of $e_{\mathrm{tr}}$, the region under the witch-hat curve is already divided into regions of lower ($P_{\mathrm{tr}}$, $\Delta e$) and higher ($P_{\mathrm{tr}}$, $\Delta e$), as observed in Fig. 4. The three selected EoSs indicate that the allowed of MMRR sequences can span categories II, III, and IV. However, the range of EoSs that can form category II TSs appears disconnected from the range of EoSs that can form the rest of the categories. In Fig. 5, we show that the NICER measurement of PSR J0740+6620 and GW170817 are responsible (from the bottom and top, respectively) for this bifurcation. Importantly, the region of EoSs right below the tip of the witch-hat curve, which majorly constructs category I TSs, is ruled out at small $e_{\mathrm{tr}}$ due to the violation of constraints from PSR J0614–3329 and GW170817. These constraints demand softer hadronic parts in the EoSs, such that the second branch satisfies them and the third branch can construct massive TSs of category I.
For intermediate value of $e_{\mathrm{tr}}$, we observe that the region of higher ($P_{\mathrm{tr}}$, $\Delta e$) is tightly constrained, which is linked with the category II TSs. In Fig. 5, we show that the mass-radius region inferred from GW170817 is mainly responsible for ruling out these EoSs. In particular, the second branches of these EoSs are too stiff (violating GW170817), followed by most of the third branches only satisfying the constraint of either of the companion masses in the GW170817 event. Instead, EoS with TSs of category III and IV are preferred in such a regime of $e_{\mathrm{tr}}$, having the third branch covering a larger mass range and satisfying the mass-radius constraints from both the companions of GW170817. The category I TSs are still ruled out in this regime of $e_{\mathrm{tr}}$ for the same reasons discussed for small $e_{\mathrm{tr}}$.
For large value of $e_{\mathrm{tr}}$, the regions of allowed EoSs shrink drastically, mainly violating constraints from PSR J0614–3329 and PSR J0348+0432. The set of allowed hadronic EoSs is rather softer, and only a few hybrid EoSs can produce massive NSs ($\gtrsim 2M_{\odot}$). Even though a larger range of $\Delta e$ is allowed in the region of higher ($P_{\mathrm{tr}}$, $\Delta e$), the range of allowed $P_{\mathrm{tr}}$ for such category II TSs is tightly constrained. Therefore, the range of allowed hadronic EoSs is heavily constrained at such a regime of $e_{\mathrm{tr}}$. Interestingly, EoSs with TSs of category IV are also unable to support massive NS and hence are strictly ruled out in this case. The TSs of category I still remain ruled out, as in previous cases; however, we observe that such TSs are allowed at extremely large $e_{\mathrm{tr}}$ (see more discussion in Sec. III.3).
Since the witch-hat curves are dictated by $c^{2}_{\mathrm{s,QM}}$ (as discussed in Sec. III.1), the allowed regions from observations will also be affected in a similar manner. In Fig. 6, we summarise the effect of $c^{2}_{\mathrm{s,QM}}$ in the picture. Since the resulting witch-hat curves are smaller with a decrease in $c^{2}_{\mathrm{s,QM}}$, the observationally constrained regions are also tighter (from right to left panel). Additionally, we inform the further inclusions (rejections) of EoSs when $3\sigma$ ($1\sigma$) confidence of observational inferences are considered. As expected, an increase in the confidence interval (from the lower to the upper panel) widens the set of allowed EoSs. Interestingly, if $3\sigma$ confidence is used to constrain EoSs with higher values of $c^{2}_{\mathrm{s,QM}}$ (0.77 and 1.0), we observe that the EoSs with category I TSs are included at large values of $e_{\mathrm{tr}}$, which remained excluded in our examination while considering $2\sigma$ confidence of observational data. Secondly, at large $e_{\mathrm{tr}}$ with $c^{2}_{\mathrm{s,QM}}=1.0$, we observe that a slim range of $P_{\mathrm{tr}}$ is allowed with $\Delta e$ being unconstrained. These hybrid EoSs construct a second branch that satisfies all the constraints while having its $P_{\mathrm{tr}}$ in the unbound region (green), as discussed in Fig. 2. However, in the interest of extending these hybrid EoSs to match the known behaviour of matter at pQCD regime (see Refs. [70, 73, 93] for such consistency treatments to agnostic EoSs), an ultimate upper bound can be drawn on $\Delta e$ ($\Delta\rho$) to be $\sim 5\times 10^{15}~\mathrm{g/cm}^{3}$ ($\sim 13\rho_{\mathrm{sat}}$).
For $1\sigma$ confidence level, hybrid EoSs are only allowed with a stiffer quark part. Additionally, the allowed EoSs specifically have lower values of $P_{\mathrm{tr}}$, indicating that only category III and IV TSs can comply with such constraints, and thereby strictly excluding category I and category II TSs. On a side note, we did not find the PSR J0030+0451 inference useful for constraining any part of this picture.
Table 1: Maximum transition energy density ($e^{\mathrm{max}}_{\mathrm{tr}}$) and respective maximum transition rest-mass density ($\rho^{\mathrm{max}}_{\mathrm{tr}}$) for various stiffnesses of quark matter EoSs using $c^{2}_{\mathrm{s,QM}}$. The values of $\Delta e$ and $\Delta\rho$ are also noted for these selected EoSs, along with their category of TSs. This computation is performed considering $2\sigma$ confidence level.
$c^{2}_{\mathrm{s,QM}}$ $e^{\mathrm{max}}_{\mathrm{tr}}~[\mathrm{g/cm}^{3}]$ $\Delta e~[\mathrm{g/cm}^{3}]$ $\rho^{\mathrm{max}}_{\mathrm{tr}}~[\rho_{\mathrm{sat}}]$ $\Delta\rho~[\rho_{\mathrm{sat}}]$ Category
1.0 $1.40\times 10^{15}$ $0.48\times 10^{15}$ $4.03$ 1.14 I
0.85 $1.27\times 10^{15}$ $0.46\times 10^{15}$ $3.70$ 1.14 I
0.55 $0.54\times 10^{15}$ $0.33\times 10^{15}$ 1.87 1.04 III
0.33 $0.33\times 10^{15}$ $0.91\times 10^{15}$ $1.18$ 2.75 II
Table 2: Maximum transition energy density ($e^{\mathrm{max}}_{\mathrm{tr}}$) and respective maximum transition rest-mass density ($\rho^{\mathrm{max}}_{\mathrm{tr}}$) under different confidence levels of constraints from observational inferences. The values of $\Delta e$ and $\Delta\rho$ are also noted for these selected EoSs, along with their category of TSs. This computation is performed considering the stiffest quark EoS ($c^{2}_{\mathrm{s,QM}}=1.0$).
Confidence $e^{\mathrm{max}}_{\mathrm{tr}}~[\mathrm{g/cm}^{3}]$ $\Delta e~[\mathrm{g/cm}^{3}]$ $\rho^{\mathrm{max}}_{\mathrm{tr}}~[\rho_{\mathrm{sat}}]$ $\Delta\rho~[\rho_{\mathrm{sat}}]$ Category
$3\sigma$ $1.40\times 10^{15}$ $0.48\times 10^{15}$ $4.03$ 1.14 I
$2\sigma$ $1.40\times 10^{15}$ $0.48\times 10^{15}$ $4.03$ 1.14 I
$1\sigma$ $0.42\times 10^{15}$ $0.49\times 10^{15}$ 1.48 1.61 III
Table 3: Maximum strength of PT discontinuity in energy density ($\Delta e_{\mathrm{max}}$) and respective rest-mass density ($\Delta\rho_{\mathrm{max}}$) for various stiffnesses of quark matter EoSs using $c^{2}_{\mathrm{s,QM}}$. The values of $e_{\mathrm{tr}}$ and $\rho_{\mathrm{tr}}$ are also noted for the allowed EoS along with their category of TSs. This computation is performed considering $2\sigma$ confidence level.
$c^{2}_{\mathrm{s,QM}}$ $\Delta e_{\mathrm{max}}~[\mathrm{g/cm}^{3}]$ $\Delta\rho_{\mathrm{max}}~[\rho_{\mathrm{sat}}]$ $e_{\mathrm{tr}}~[\mathrm{g/cm}^{3}]$ $\rho_{\mathrm{tr}}~[\rho_{\mathrm{sat}}]$ Category
1.0 $2.94\times 10^{15}$ 7.76 $0.90\times 10^{15}$ 2.92 II
0.85 $2.35\times 10^{15}$ 6.29 $0.88\times 10^{15}$ 2.85 II
0.55 $1.43\times 10^{15}$ 4.10 $0.49\times 10^{15}$ 1.67 II
0.33 $0.92\times 10^{15}$ 2.80 $0.32\times 10^{15}$ 1.12 II
Table 4: Maximum strength of PT discontinuity in energy density ($\Delta e_{\mathrm{max}}$) and respective rest-mass density ($\Delta\rho_{\mathrm{max}}$) under different confidence levels of constraints from observational inferences. The values of $e_{\mathrm{tr}}$ and $\rho_{\mathrm{tr}}$ are also noted for the allowed EoS along with their category of TSs. This computation is done for the stiffest quark EoS ($c^{2}_{\mathrm{s,QM}}=1.0$).
Confidence $\Delta e_{\mathrm{max}}~[\mathrm{g/cm}^{3}]$ $\Delta\rho_{\mathrm{max}}~[\rho_{\mathrm{sat}}]$ $e_{\mathrm{tr}}~[\mathrm{g/cm}^{3}]$ $\rho_{\mathrm{tr}}~[\rho_{\mathrm{sat}}]$ Category
$3\sigma$ $4.97\times 10^{15}$ 13.13 $0.90\times 10^{15}$ 2.92 II
$2\sigma$ $2.94\times 10^{15}$ 7.76 $0.90\times 10^{15}$ 2.92 II
$1\sigma$ $0.68\times 10^{15}$ 2.21 $0.30\times 10^{15}$ 1.06 III
Table 5: Maximum mass of TSs ($M^{\mathrm{max}}_{\mathrm{TS}}$) and their categories, for various stiffnesses of quark matter EoSs using $c^{2}_{\mathrm{s,QM}}$. The values of $\Delta e$ and $\Delta\rho$ are also noted for the allowed EoS that construct the stellar configurations supporting $M^{\mathrm{max}}_{\mathrm{TS}}$. This computation is performed considering $2\sigma$ confidence level.
$c^{2}_{\mathrm{s,QM}}$ $M^{\mathrm{max}}_{\mathrm{TS}}~[M_{\odot}]$ $e_{\mathrm{tr}}~[\mathrm{g/cm}^{3}]$ $\Delta e~[\mathrm{g/cm}^{3}]$ $\rho_{\mathrm{tr}}~[\rho_{\mathrm{sat}}]$ $\Delta\rho~[\rho_{\mathrm{sat}}]$ Category
1.0 2.05 $1.26\times 10^{15}$ $0.45\times 10^{15}$ $3.66$ 1.10 I
0.85 2.02 $1.21\times 10^{15}$ $0.45\times 10^{15}$ $3.55$ 1.10 I
0.55 1.57 $0.30\times 10^{15}$ $1.09\times 10^{15}$ $1.03$ 3.22 II
0.33 1.38 $0.30\times 10^{15}$ $0.67\times 10^{15}$ $1.03$ 2.17 II
Table 6: Maximum mass of TSs ($M^{\mathrm{max}}_{\mathrm{TS}}$) and their categories, under different confidence levels of constraints from observational inferences. The values of $\Delta e$ and $\Delta\rho$ are also noted for the allowed EoS that construct the stellar configurations supporting $M^{\mathrm{max}}_{\mathrm{TS}}$. This computation is performed considering the stiffest quark EoS ($c^{2}_{\mathrm{s,QM}}=1.0$).
Confidence $M^{\mathrm{max}}_{\mathrm{TS}}~[M_{\odot}]$ $e_{\mathrm{tr}}~[\mathrm{g/cm}^{3}]$ $\Delta e~[\mathrm{g/cm}^{3}]$ $\rho_{\mathrm{tr}}~[\rho_{\mathrm{sat}}]$ $\Delta\rho~[\rho_{\mathrm{sat}}]$ Category
$3\sigma$ 2.32 $0.97\times 10^{15}$ $0.39\times 10^{15}$ 2.92 0.94 I
$2\sigma$ 2.05 $1.26\times 10^{15}$ $0.45\times 10^{15}$ 3.66 1.01 I
$1\sigma$ 1.37 $0.3\times 10^{15}$ $0.67\times 10^{15}$ 1.03 2.17 III
III.3 Limits on properties of PT and maximum mass of TSs
In Table 1, we note the EoSs with the maximum value of $e_{\mathrm{tr}}$ ($e_{\mathrm{tr}}^{\mathrm{max}}$), for varying stiffness of quark EoS. Interestingly, the EoSs that form category I TSs are allowed at extremely large $e_{\mathrm{tr}}$ ($\approx 4\rho_{\mathrm{sat}}$), and are consequently responsible for setting the upper bound on $e_{\mathrm{tr}}$. Such hybrid EoSs are soft enough to construct second branches that follow all the constraints ($2\sigma$ confidence level) while supporting massive TSs on the third branch. Since $P_{\mathrm{tr}}$ of these EoSs lie in the bound region (blue) of the witch-hat curve, the maximum of $\Delta e$ is strictly decided from TOV solutions. In the cases of softer quark EoSs, the hybrid EoSs do not support massive TSs to satisfy the mass constraint from PSR J0348+0432. It can be anticipated from Fig. 3 where we show that smaller $c^{2}_{\mathrm{s,QM}}$ sets smaller $P_{\mathrm{tr}}^{\mathrm{max}}$. It suggests that a second branch supports massive NSs ($\gtrsim 2.0~M_{\odot}$) that will have central pressures higher than $P_{\mathrm{tr}}^{\mathrm{max}}$, and for such EoSs, TS solutions would not exist. In such cases, $e_{\mathrm{tr}}^{\mathrm{max}}$ is achieved via EoSs with TSs of categories II–III, which is the reason behind the $e_{\mathrm{tr}}^{\mathrm{max}}$ being drastically smaller than we found for stiffer quark EoSs. In Table 2, we find the $e_{\mathrm{tr}}^{\mathrm{max}}$ for various confidence intervals using the stiffest quark EoS. The results remain unchanged for $2\sigma$ and $3\sigma$ confidence levels, as these results are mainly bound by mass constraint from PSR J0348+0432. For $1\sigma$, the observational constraints demand even more compact second branch around the canonical NS mass ($1.4~M_{\odot}$). A hybrid EoS with such a soft hadronic part fails to support massive TSs of category I. In compliance with $1\sigma$ confidence level, $e_{\mathrm{tr}}^{\mathrm{max}}$ is established at a much smaller value by an EoS with TSs of category III.
Similarly, in Tables 3 and 4, we note the EoSs with the maximum value of $\Delta e$ ($\Delta e_{\mathrm{max}}$). These inferences are made out of EoSs that construct TSs of category II (except for the case of $1\sigma$ confidence interval that only allows TSs of category III and IV). For a stiffer quark EoS, the hybrid EoSs construct a second branch that satisfies all observations except for PSR J0614–3329. The third branch of category II kind then satisfies PSR J0614–3329, resulting in inclusion of such an EoS with large $\Delta e$. For a softer quark EoS, the hybrid EoSs construct a second branch that satisfies the observation PSR J0740+6630, whereas the third branch of category II kind then satisfies the rest of the constraints. For $1\sigma$ confidence level, an EoS with TSs of category III decides the $\Delta e_{\mathrm{max}}$, where the third branch satisfies all the constraints.
In Table 5, we noted the maximum mass of the TSs ($M^{\mathrm{max}}_{\mathrm{TS}}$) considering various stiffnesses of quark matter EoS (not to be confused with $M_{\mathrm{TOV,3}}$ as defined in Sec. II.1). For soft quark EoSs, we observe that the $M^{\mathrm{max}}_{\mathrm{TS}}$ is achieved for $\rho_{\mathrm{tr}}\sim 1\rho_{\mathrm{sat}}$, which is at the limit of our analysis. Higher $M^{\mathrm{max}}_{\mathrm{TS}}$ can be found for lower $\rho_{\mathrm{tr}}$, but it will be very low for a hadron-to-quark PT. In Table 6, we noted the $M^{\mathrm{max}}_{\mathrm{TS}}$ considering various confidence intervals of observational constraints. For $2\sigma$ and $3\sigma$ confidence, we observe that $M^{\mathrm{max}}_{\mathrm{TS}}>2.0~M_{\odot}$, constructed out of category I TSs. For $1\sigma$ confidence level infers $M^{\mathrm{max}}_{\mathrm{TS}}=1.37~M_{\odot}$, which is drastically smaller, and is a manifestation of allowing only category III and IV TSs.
IV Conclusions
We performed a thorough analysis of hybrid EoSs with strong first-order (Maxwell-type) PT, which can form TS solutions. We employ an agnostic approach for the systematic construction of EoSs mainly using the four parameters—transition energy density ($e_{\mathrm{tr}}$) and transition pressure ($P_{\mathrm{tr}}$), the strength of discontinuity in energy density ($\Delta e$) and constant speed of sound for quark matter EoS ($c^{2}_{\mathrm{s,QM}}$).
Given the setup of our parameter space, we systematically present a comprehensive picture of the conditions required by the hybrid EoSs to form TSs. These conditions are purely achieved from TOV solutions within the causality limit. For a fixed value of $e_{\mathrm{tr}}$ and $c^{2}_{\mathrm{s,QM}}$, we define three regions of $P_{\mathrm{tr}}$:
  • solution-less region–TSs do not form for any $\Delta e$
  • bound region–TSs form for a range of $\Delta e$
  • unbound region–TSs form beyond a minimum limit on $\Delta e$
In this article, the loci of minimum and maximum $\Delta e$ are tossed as the ‘witch-hat’ curves. We show that a witch-hat curve can be described for a given ($e_{\mathrm{tr}}$, $c^{2}_{\mathrm{s,QM}}$), where each point ($P_{\mathrm{tr}}$, $\Delta e$) under this curve represents a hybrid EoS that can form TSs. We find that an increase in the values of $e_{\mathrm{tr}}$ or $c^{2}_{\mathrm{s,QM}}$ enlarges its witch-hat curve, thereby allowing a larger range of EoSs to form TSs. For certain combinations of ($e_{\mathrm{tr}}$, $c^{2}_{\mathrm{s,QM}}$), witch-hat curves can be punctured (or incomplete) based on the causality limit. We observed that the TOV solutions can put only an upper limit on $P_{\mathrm{tr}}$ (if not by the causality limit). No upper limit is provided for the $e_{\mathrm{tr}}$ or $\Delta e$ (at small values of $P_{\mathrm{tr}}$). However, the recent observational measurements provide strong constraints on this picture.
We explain the constrained picture from the observational inferences ($2\sigma$ confidence level) on the region under the witch-hat curves for three values of $e_{\mathrm{tr}}$ (qualitatively at small, intermediate and large values) with fixed $c^{2}_{\mathrm{s,QM}}=1.0$ (stiffest quark EoS). We find that different observational measurements are relevant for constraining different regions under the witch-hat curves, and that different combinations of measurements come into play at different regimes of $e_{\mathrm{tr}}$. The constraints split the allowed region under the witch-hat curve into two fractions—($i$) region of lower ($P_{\mathrm{tr}}$, $\Delta e$), forming TSs of category III and IV, and ($ii$) region of higher ($P_{\mathrm{tr}}$, $\Delta e$), forming TSs of category II. We observe that TSs of category I are only allowed at extremely large values of $e_{\mathrm{tr}}$ ($\approx 4\rho_{\mathrm{sat}}$). As the $e_{\mathrm{tr}}$ increases, the observational inferences constrain the regions under the witch-hat curves more tightly.
Importantly, a maximum limit on $\Delta e$ is sufficiently provided for all values of $e_{\mathrm{tr}}$. We note that a maximum limit on $e_{\mathrm{tr}}$ is also provided, as an increase in $e_{\mathrm{tr}}$ allows softer hadronic EoSs which are tightly constrained when such hybrid EoSs are unable to support massive NSs ($\gtrsim 2.0~M_{\odot}$). On the other hand, we found a maximum mass for the TSs. While keeping the stiffest quark EoS ($c^{2}_{\mathrm{s,QM}}=1.0$), we report the following thermodynamic and astrophysical upper bounds:
  • $e^{\mathrm{max}}_{\mathrm{tr}}=1.4\times 10^{15}~\mathrm{g/cm}^{3}$ ($\rho^{\mathrm{max}}_{\mathrm{tr}}=4.03\rho_{\mathrm{sat}}$), governed by an EoS with TSs of category I
  • $\Delta e_{\mathrm{max}}=2.94\times 10^{15}~\mathrm{g/cm}^{3}$ ($\Delta\rho_{\mathrm{max}}=7.75\rho_{\mathrm{sat}}$), governed by an EoS with TSs of category II
  • $M^{\mathrm{max}}_{\mathrm{TS}}=2.05~M_{\odot}$, governed by an EoS with TSs of category I
Naturally, for a softer quark EoS, these limits will be affected, the results of which are summarised in the dedicated tables. We note that the contributing categories of TSs change when softer quark EoSs are considered.
[…]
Observation of TSs (especially massive ones like the $M^{\mathrm{max}}_{\mathrm{TS}}$), can be smoking gun signatures of PT in the density regimes present in the cores of NSs. As found in this study, these can immediately place strong constraints on the $e_{\mathrm{tr}}$ and $\Delta e$, and significantly further our understanding of the possible nature of hadron-to-quark PT that can exist at such high densities. It would be interesting to consider additional effects, such as rotation and magnetic field, to gain a complete insight into this picture. A complementary study should also be drawn by considering a different class of PT with a non-vanishing speed of sound (for instance, the impact of varying surface tension at the interface of phases [138, 139, 140, 141, 142, 143]). A parallel study has implemented speed of sound parametrisation to define the hadronic and quark EoSs for the class of EoSs that construct TS [93], in contrast to the implementation of a piecewise polytrope with the CSS formalism used in this study.
Acknowledgements
It is a pleasure to thank Sagnik Chatterjee for his crucial insights into this work. The authors also express their gratitude to the Deucalion HPC in Portugal for its support in the Advanced Computing Project 2025.00067.CPCA A3, RNCA (Rede Nacional de Computação Avançada), financed by the FCT (Fundação para a Ciência e a Tecnologia, IP, Portugal). The test setups were completed in the National Supercomputing Mission (NSM) resource ‘PARAM Ganga’ at IIT Roorkee, which is implemented by C-DAC and supported by the Ministry of Electronics and Information Technology (MeitY) and the Department of Science and Technology (DST), Government of India. The small-scale tests were conducted at the Bhaskara HPC in IISER Bhopal. SH and RM acknowledge the SERB grant: CRG/2022/000663. This research work is done using the packages—Numpy [144], SciPy [145], Matplotlib [146], Seaborn [147], Jupyter [148], Numba [149] and OpenCV [150].